OpenStudy (anonymous):

Solve: 3x-5y=17 9x-15y=-4

OpenStudy (solomonzelman):

multiply the first equation times 3, and tell me what you think about it.

OpenStudy (anonymous):

I got 9x-15y=51

OpenStudy (solomonzelman):

yes. So in the second equation the "9x-15y" is -4, and in the second equation it is 51.

OpenStudy (anonymous):

Wouldn't you multiply the second equation by -1? since you have to have opposite coefficients.

OpenStudy (solomonzelman):

9x-15y=51 9x-15y=-4 see that the "9x-15y" is equal to different numbers?

OpenStudy (solomonzelman):

in first equation it is 51, and in a second equation it is -4.

OpenStudy (solomonzelman):

Therefore there are no solutions because 51 is not equal to -4.

OpenStudy (anonymous):

ok thanks